# Mathematical Language Models: A Survey

### 14 Dec 2023 v2

@@ 4.4 Advanced CoT Methods
Furthermore, the self-refine approach, as proposed by Madaan et al. [madaan2023self] and Sun et al. [sun2023adaplanner], focuses on error correction and summarizing past experiences. Specifically, Self-Refine performs pre-execution validation, providing feedback to the model for refinement in case of errors. Zhou et al. [zhou2023isr] then took a step further by merging the iterative self-refine with PDDL in ISR-LLM. Additionally, Shinn et al. [shinn2023reflexion] introduced re-flexion to rectify errors stemming from previous actions.
@@
The Socratic teaching method [nelson1980socratic], derived from the philosophies of ancient Greek thinker Socrates, provides a unique perspective when considering the integration of LMs in mathematics education [davies2021advancing]. This method, emphasizing questioning and dialogue, aims to guide learners toward profound thinking [davis1997dialogue, brickhouse2009socratic]. In realm of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery.
For LLMs, research shows that the Socratic teaching method can also improve their abilities to a certain extent. For example, the framework of Socratic Models (SMs) [zeng2022socratic] significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. In Chang et al. [chang2023prompting], techniques including definition, dialectics, cross-examination, induction, and counterfactual reasoning are utilized. These techniques contribute to enhancing the LLMs’ comprehension of user intentions, thereby improving model’s effectiveness in addressing user queries. The inductive, deductive, and analogical reasoning techniques in the Socratic teaching method establish a comprehensive analytical framework for solving mathematical problems [chang2023prompting, alhossami2023language], ensuring a strong and concise ability to address the complexities of mathematical problem-solving. Moreover, contributions from Ang et al. [ang2023socratic] and Al et al. [al2023socratic] provided invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability.
Overall, the integration of Socratic teaching techniques in LLMs offers novel perspectives and tools for mathematics education. Unlike CoT [wei2022chain], Socratic questioning explicitly guides the thinking process, stimulating effective recursive thinking and displaying greater resilience to errors in the thinking process [qi2023art]. Ranging from basic conversational guidance to intricate problem-solving, these methods demonstrate the effective fusion of modern technology with classical ideas to bolster the learning and comprehension abilities of LLMs.
@@
VERBPHYSICS [forbes2017verb] 733 1,096 1,828 3,657 EN Calculation Training Formula
Clinical [spithourakis2016numerically, spithourakis2018numeracy] 11,170 1,625 3,220 16,015 EN Calculation Training Formula
Scientific [spithourakis2018numeracy] 14,694 2,037 4,231 20,962 EN Calculation Training Formula
DoQ [elazar2019large] 587 5,418 6,007 12,012 EN Calculation Training Formula
DROP [dua2019drop] 77,409 9,536 9,622 96,567 EN Calculation Training Text
AddSub [hosseini-etal-2014-learning] - - - 395 EN MWP Training Formula
SingleOp [roy2015reasoning] 265 107 159 531 EN MWP Training Formula
SingleEq [koncel2015parsing] - - - 508 EN MWP Training Formula
MultiArith [roy2015solving] 420 - 180 600 EN MWP Training Formula
Alg514 [kushman-etal-2014-learning] - - - 514 EN MWP Training Formula
Math23k [wang2017deep] 22,162 - 1000 23,162 CH MWP Training Formula
AQuA [ling2017program] 97,467 254 254 97,975 EN MWP Training Text
MathQA [amini2019mathqa] 29,837 4,475 28,985 37,297 EN MWP Training Formula
GSM8K [cobbe2021training] 7,473 1,319 8,792 EN MWP Training Text
SVAMP [patel2021nlp] 700 300 1,000 EN MWP Training Formula
DRAW [upadhyay2015draw] - - - 1,000 EN MWP Training Formula
Dolphin1878 [shi2015automatically] - 374 1,504 1,878 EN MWP Training Formula
AllArith [roy2017unit] - - - 831 EN MWP Training Formula
DRAW-1K [upadhyay2017annotating] - - - 1,000 EN MWP Training Formula
HMWP [qin2020semantically] - - - 5,470 CH MWP Training Formula
ArMATH [alghamdi2022armath] - - - 6,000 Arabic MWP Training Formula
TabMWP [lu2022dynamic] - - - 38,431 EN MWP Training Text
Ape210K [zhao2020ape210k] - - - 210,488 CH MWP Training Formula
@@
FinQA [chen2021finqa] 6,251 883 1,147 8,281 EN MWP Training Formula
REALFP [kalyan2021much] 185 185 558 928 EN MWP Training Formula
SYNTHFP [kalyan2021much] 10,000 1,000 1,000 12,000 EN MWP Training Formula
TAT-QA [zhu2021tat] - - - 16,552 EN MWP Training Text
MultiHiertt [zhao2022multihiertt] 7,830 1,044 1,566 10,440 EN MWP Training Formula
MML [grabowski2015four] - - - 57,882 EN TP Training Formula
HolStep [kaliszyk2016holstep] 2,013,046 - 196,030 2,209,076 EN TP Training Formula
Feit-Thompson [huang2018gamepad] - - - 83,478 EN TP Training Formula
CoqGym [yang2019learning] - - - 71,000 EN TP Training Formula
HOList [bansal2019holist] - - - 29,462 EN TP Training Formula
IsarStep [li2020isarstep] 820,000 5,000 5,000 830,000 EN TP Training Formula
LISA [jiang2021lisa] - - - 183,000 EN TP Training Formula
NaturalProofs [welleck2021naturalproofs] 32,000 EN TP Training Text
LeanStep [han2021proof] - - - 21,606,000 EN TP Training Formula
NumGLUE [mishra2022numglue] - - - 101,835 EN Calculation Benchmark Text
Dophin18k [huang2016well] - - - 18,460 EN MWP Benchmark Text
MAWPS [koncel2016mawps] - - - 3,320 EN MWP Benchmark Formula
ASDiv [miao2020diverse] - - - 2,305 EN MWP Benchmark Formula
MATH [hendrycks2021measuring] 7,500 5,000 12,500 EN MWP Benchmark Text
MGSM [shi2022language] - - - - Multilingual MWP Benchmark Text
Mathematics [saxton2018analysing] 2,000,000 100,000 2,100,000 EN MWP Benchmark Formula
MMLU-Math [hendryckstest2021] - - - 906 EN WMP Benchmark Formula
AGIEval [zhong2023agieval] - - - 469/220 CH/EN MWP Benchmark Formula
MATH 401 [yuan2023well] - - - 401 EN MWP Benchmark Formula
INT [wu2020int] - - - - EN TP Benchmark Formula
miniF2F [zheng2021minif2f] - 244 244 488 EN TP Benchmark Formula
Aggregate [roy2018mapping] - - - 1,492 EN MWP Augmented Formula
MathQA-Python [austin2021program] 19,209 2,822 2,822 23,914 EN MWP Augmented Code
Math50k [li2023camel] - - - 50,000 EN WMP Augmented Text
PRM800K [lightman2023lets] - - - 2,868 EN WMP Augmented Text
MetaMathQA [yu2023metamath] - - - 395,000 EN MWP Augmented Text
Lila [mishra2022lila] - - - 134,000 EN MWP Augmented Code
PEN [kim2022ept] - - - 3,581 EN MWP Augmented Formula
miniF2F+informal [jiang2022draft] - 244 244 488 EN TP Augmented Formula
NaturalProofs-Gen [welleck2022naturalprover] 12,500 1,000 1,000 14,500 EN TP Augmented Text
@@
Table 1: The statistics information of mathematical datasets. Solution means the format of the output, such as text, formula and code.
5 Datasets
@@
To train and evaluate the arithmetic and math ability of language models, various math word problems datasets [saxton2018analysing, hendrycks2021measuring, cobbe2021training, kushman-etal-2014-learning] are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1).
5.1 Training Datasets
5.1.1 Mathematical Calculation
@@
Several studies have introduced datasets aiming to identify numerical information within text [spithourakis2016numerically, spithourakis2018numeracy, elazar2019large], specifically targeting the prediction of attribute values. The Clinical Data [spithourakis2016numerically, spithourakis2018numeracy] consists of clinical records sourced from the London Chest Hospital. Each patient record includes a text report and structured KB tuples detailing 20 potential numeric attributes like age and gender. Scientific Data [spithourakis2018numeracy] encompasses paragraphs extracted from Cornell’s ARXIV repository, featuring over half a million converted papers across 37 scientific sub-fields. Additionally, Elazar et al. [elazar2019large] proposed the Distributions over Quantities (DoQ) dataset, comprising empirical counts of scalar attribute values linked to more than 350K nouns, adjectives, and verbs across 10 different attributes.
@@
Furthermore, certain studies consider the relational knowledge between various numbers. For instance, the VERBPHYSICS [forbes2017verb] dataset aggregates crowdsourced information about actions and objects, encompassing the relative knowledge of grounded object pairs and implications of actions associated with those objects. More intricately, the DROP [dua2019drop] dataset necessitates Discrete Reasoning Over the content of Paragraphs. In this benchmark of 55,000 adversarially-created questions, a system is tasked with resolving references within a question, potentially spanning multiple input positions, and performing discrete operations like addition, counting, or sorting.
5.1.2 Math Word Problems
@@
A large number of datasets are proposed for MWPs. Hosseini et al. [hosseini-etal-2014-learning] curated the AddSub dataset, primarily focusing on addition and subtraction problems (Figure 9). This dataset identifies relevant variables and their respective values to translate sentences into problem statements, represented as equations. Another dataset, SingleOp [roy2015reasoning], encompasses a wider range of mathematical operations including multiplication and division. Its purpose is to facilitate reasoning about quantities articulated in natural language. DRAW [upadhyay2015draw] comprises 1000 algebra word problems semi-automatically annotated for evaluating automatic solvers. It features gold coefficient alignments crucial for uniquely identifying equation system derivations. The Alg514 [kushman-etal-2014-learning] dataset comprises 514 linear algebra problems structured around 28 equation templates. Meanwhile, MultiArith [roy2015solving] is designed to handle arithmetic problems involving multiple steps and operations, without predefined templates.
@@
Several studies have introduced methods focused on predicting equations through semantic parsing. The SingleEq [koncel2015parsing] dataset revolves around grade-school algebra word problems. It comprises samples that can be correlated to a single equation involving multiple mathematical operations on one variable, structured as a parsing tree (Figure 10). In contrast, the Dolphin1878 [shi2015automatically] dataset contains 1,878 word problems in mathematics, wherein each sample might encompass multiple equations and variables depicted through multiple trees. AllArith [roy2017unit] offers a concise representation of dependencies among number units mentioned in a given problem, referred to as Unit Dependency Graphs (UDGs). To augment dataset size, Math23k [wang2017deep] gathers 23,161 problems labeled with structured equations and corresponding answers. GSM8K [cobbe2021training] comprises 8.5K high-quality grade school math problems crafted by human problem writers. These problems typically involve 2 to 8 steps for resolution, predominantly requiring a sequence of elementary calculations using basic arithmetic operations to arrive at the final solution.
Figure 9: An example of AddSub, taken from [hosseini-etal-2014-learning].
Figure 10: An example of SingleEQ, taken from [roy2015reasoning].
@@
HMWP [qin2020semantically] aims to enhance dataset diversity by extracting math word problems from a Chinese K12 problem bank. This effort sought to validate the universality of math word problem solvers and expand research in MWPs to better align with real-world scenarios. The dataset encompasses three types of MWPs: arithmetic word problems, equation set problems, and non-linear equation problems. In total, it comprises 5,491 MWPs, including 2,955 single-variable linear MWPs, 1,636 two-variable linear MWPs, and 900 single-variable nonlinear MWPs. Additionally, MathQA [amini2019mathqa] presents a diverse collection of 37,000 English multiple-choice math word problems spanning various mathematical domains (Figure 11).
Figure 11: An example of MathQA, taken from [amini2019mathqa].
@@
Existing research endeavors to enhance the comprehensibility and precision of intricate reasoning by providing fine-grained annotations of reasoning processes [upadhyay2017annotating, ling2017program, chen2021finqa, lightman2023lets, zhao2022multihiertt]. DRAW-1K [upadhyay2017annotating] introduces a novel dataset featuring 1,000 general algebra word problems, each annotated with derivations. AQUA [ling2017program] structures each question into four components: problem description, multiple-choice answer options, rationale, and correct option label (Figure 12). FinQA [chen2021finqa] presents an expert-annotated dataset comprising 8,281 financial QA pairs, meticulously annotated with numerical reasoning processes to ensure comprehensive explication. PRM800K [lightman2023lets] assigns labels (positive, negative, or neutral) to each step in solving MATH problems sampled by a large-scale generator. The training set encompasses 800K step-level labels across 75K solutions to 12K problems.
Figure 12: An example of AQUA, taken from [ling2017program].
@@
Additionally, Patel et al. [patel2021nlp] introduced the SVAMP challenge set to establish a more robust evaluation framework for methods designed to tackle elementary-level math word problems. Alghamdi et al. [alghamdi2022armath] contributed the inaugural large-scale dataset, ArMATH, for Arabic MWPs. This dataset comprises 6,000 samples of primary-school math problems written in Modern Standard Arabic. Kalyan et al. [kalyan2021much] proposed two datasets, REALFP and SYNTHFP, for a novel task called Fermi Problems (FPs). FPs entail questions whose answers can only be approximately estimated due to the impracticality or impossibility of precise computation. This task necessitates the amalgamation of various reasoning abilities, including suitable abstractions, commonsense knowledge, and the creative synthesis of problem-solving strategies.
@@
Recently, mathematical reasoning has incorporated multiple input modalities such as textual and tabular data [lu2022dynamic, zhu2021tat, zhao2022multihiertt]. For instance, TabMWP [lu2022dynamic] presents a new dataset comprising 38,431 open-domain grade-level problems that necessitate mathematical reasoning across both textual and tabular data. Zhu et al. [zhu2021tat] introduced a large-scale QA dataset, TAT-QA, encompassing both tabular and textual data. This dataset often requires numerical reasoning—such as addition, subtraction, multiplication, division, counting, comparison/sorting, and their combinations—based on real financial reports to infer answers. Additionally, MultiHiertt [zhao2022multihiertt] is a proposed dataset consisting of QA pairs derived from Multi Hierarchical Tabular and Textual data (Figure 13). Each document in this dataset contains multiple tables and longer unstructured texts, accompanied by fine-grained annotations of reasoning processes and supporting facts to elucidate complex numerical reasoning.
Figure 13: An example of MultiHiertt, taken from [zhao2022multihiertt].
5.1.3 Theorem Proving

### 23 Feb 2024 v3

@@ 4.4. Advanced CoT Methods
Furthermore, the self-refine approach, as proposed by Madaan et al. (Madaan et al., 2023) and Sun et al. (Sun et al., 2023), focuses on error correction and summarizing past experiences. Specifically, Self-Refine performs pre-execution validation, providing feedback to the model for refinement in case of errors. Zhou et al. (Zhou et al., 2023b) then took a step further by merging the iterative self-refine with PDDL in ISR-LLM. Additionally, Shinn et al. (Shinn et al., 2023) introduced re-flexion to rectify errors stemming from previous actions.
@@
The Socratic teaching method (Nelson, 1980), derived from the philosophies of ancient Greek thinker Socrates, provides a unique perspective when considering the integration of LMs in mathematics education (Davies et al., 2021). This method, emphasizing questioning and dialogue, aims to guide learners toward profound thinking (Davis and Steinglass, 1997; Brickhouse and Smith, 2009). In realm of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery.
For LLMs, research shows that the Socratic teaching method can also improve their abilities to a certain extent. For example, the framework of Socratic Models (SMs) (Zeng et al., 2022) significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. In Chang et al. (Chang, 2023), techniques including definition, dialectics, cross-examination, induction, and counterfactual reasoning are utilized. These techniques contribute to enhancing the LLMs’ comprehension of user intentions, thereby improving model’s effectiveness in addressing user queries. The inductive, deductive, and analogical reasoning techniques in the Socratic teaching method establish a comprehensive analytical framework for solving mathematical problems (Chang, 2023; Al-Hossami et al., 2023a), ensuring a strong and concise ability to address the complexities of mathematical problem-solving. Moreover, contributions from Ang et al. (Ang et al., 2023) and Al et al. (Al-Hossami et al., 2023b) provided invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability.
Overall, the integration of Socratic teaching techniques in LLMs offers novel perspectives and tools for mathematics education. Unlike CoT (Wei et al., 2022), Socratic questioning explicitly guides the thinking process, stimulating effective recursive thinking and displaying greater resilience to errors in the thinking process (Qi et al., 2023). Ranging from basic conversational guidance to intricate problem-solving, these methods demonstrate the effective fusion of modern technology with classical ideas to bolster the learning and comprehension abilities of LLMs.
@@
VERBPHYSICS (Forbes and Choi, 2017) 733 1,096 1,828 3,657 EN Calculation Training Formula
Clinical (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) 11,170 1,625 3,220 16,015 EN Calculation Training Formula
Scientific (Spithourakis and Riedel, 2018) 14,694 2,037 4,231 20,962 EN Calculation Training Formula
DoQ (Elazar et al., 2019) 587 5,418 6,007 12,012 EN Calculation Training Formula
DROP (Dua et al., 2019) 77,409 9,536 9,622 96,567 EN Calculation Training Text
AddSub (Hosseini et al., 2014) - - - 395 EN MWP Training Formula
SingleOp (Roy et al., 2015) 265 107 159 531 EN MWP Training Formula
SingleEq (Koncel-Kedziorski et al., 2015) - - - 508 EN MWP Training Formula
MultiArith (Roy and Roth, 2015) 420 - 180 600 EN MWP Training Formula
Alg514 (Kushman et al., 2014) - - - 514 EN MWP Training Formula
Math23k (Wang et al., 2017) 22,162 - 1000 23,162 CH MWP Training Formula
AQuA (Ling et al., 2017) 97,467 254 254 97,975 EN MWP Training Text
MathQA (Amini et al., 2019) 29,837 4,475 28,985 37,297 EN MWP Training Formula
GSM8K (Cobbe et al., 2021) 7,473 1,319 8,792 EN MWP Training Text
SVAMP (Patel et al., 2021) 700 300 1,000 EN MWP Training Formula
DRAW (Upadhyay and Chang, 2015) - - - 1,000 EN MWP Training Formula
Dolphin1878 (Shi et al., 2015) - 374 1,504 1,878 EN MWP Training Formula
AllArith (Roy and Roth, 2017) - - - 831 EN MWP Training Formula
DRAW-1K (Upadhyay and Chang, 2017) - - - 1,000 EN MWP Training Formula
HMWP (Qin et al., 2020) - - - 5,470 CH MWP Training Formula
ArMATH (Alghamdi et al., 2022) - - - 6,000 Arabic MWP Training Formula
TabMWP (Lu et al., 2023b) - - - 38,431 EN MWP Training Text
Ape210K (Zhao et al., 2020) - - - 210,488 CH MWP Training Formula
@@
FinQA (Chen et al., 2021) 6,251 883 1,147 8,281 EN MWP Training Formula
REALFP (Kalyan et al., 2021) 185 185 558 928 EN MWP Training Formula
SYNTHFP (Kalyan et al., 2021) 10,000 1,000 1,000 12,000 EN MWP Training Formula
TAT-QA (Zhu et al., 2021) - - - 16,552 EN MWP Training Text
MultiHiertt (Zhao et al., 2022) 7,830 1,044 1,566 10,440 EN MWP Training Formula
MATHPILE (Wang et al., 2023c) - - - 903,180 EN MWP Training Text
OpenWebMath (Paster et al., 2023) - - - - EN MWP Training Formula
MML (Grabowski et al., 2015) - - - 57,882 EN TP Training Formula
HolStep (Kaliszyk et al., 2017) 2,013,046 - 196,030 2,209,076 EN TP Training Formula
Feit-Thompson (Huang et al., 2019) - - - 83,478 EN TP Training Formula
CoqGym (Yang and Deng, 2019) - - - 71,000 EN TP Training Formula
HOList (Bansal et al., 2019) - - - 29,462 EN TP Training Formula
IsarStep (Li et al., 2021) 820,000 5,000 5,000 830,000 EN TP Training Formula
LISA (Jiang et al., 2021) - - - 183,000 EN TP Training Formula
NaturalProofs (Welleck et al., 2021) 32,000 EN TP Training Text
LeanStep (Han et al., 2022) - - - 21,606,000 EN TP Training Formula
NumGLUE (Mishra et al., 2022b) - - - 101,835 EN Calculation Benchmark Text
Dophin18k (Huang et al., 2016) - - - 18,460 EN MWP Benchmark Text
MAWPS (Koncel-Kedziorski et al., 2016) - - - 3,320 EN MWP Benchmark Formula
ASDiv (Miao et al., 2020) - - - 2,305 EN MWP Benchmark Formula
MATH (Hendrycks et al., 2021b) 7,500 5,000 12,500 EN MWP Benchmark Text
MGSM (Shi et al., 2022) - - - - Multilingual MWP Benchmark Text
Mathematics (Saxton et al., 2019) 2,000,000 100,000 2,100,000 EN MWP Benchmark Formula
MMLU-Math (Hendrycks et al., 2021a) - - - 906 EN WMP Benchmark Formula
AGIEval (Zhong et al., 2023) - - - 469/220 CH/EN MWP Benchmark Formula
MATH 401 (Yuan et al., 2023b) - - - 401 EN MWP Benchmark Formula
INT (Wu et al., 2021) - - - - EN TP Benchmark Formula
miniF2F (Zheng et al., 2023) - 244 244 488 EN TP Benchmark Formula
Aggregate (Roy and Roth, 2018) - - - 1,492 EN MWP Augmented Formula
MathQA-Python (Austin et al., 2021) 19,209 2,822 2,822 23,914 EN MWP Augmented Code
Math50k (Li et al., 2023a) - - - 50,000 EN WMP Augmented Text
PRM800K (Lightman et al., 2023) - - - 2,868 EN WMP Augmented Text
MetaMathQA (Yu et al., 2023b) - - - 395,000 EN MWP Augmented Text
Lila (Mishra et al., 2022a) - - - 134,000 EN MWP Augmented Code
PEN (Kim et al., 2022) - - - 3,581 EN MWP Augmented Formula
miniF2F+informal (Jiang et al., 2023b) - 244 244 488 EN TP Augmented Formula
NaturalProofs-Gen (Welleck et al., 2022) 12,500 1,000 1,000 14,500 EN TP Augmented Text
@@
Table 1. The statistics information of mathematical datasets. Solution means the format of the output, such as text, formula and code.
5. Datasets
@@
To train and evaluate the arithmetic and math ability of language models, various math word problems datasets (Saxton et al., 2019; Hendrycks et al., 2021b; Cobbe et al., 2021; Kushman et al., 2014) are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1).
5.1. Training Datasets
5.1.1. Mathematical Calculation
@@
Several studies have introduced datasets aiming to identify numerical information within text (Spithourakis et al., 2016; Spithourakis and Riedel, 2018; Elazar et al., 2019), specifically targeting the prediction of attribute values. The Clinical Data (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) consists of clinical records sourced from the London Chest Hospital. Each patient record includes a text report and structured KB tuples detailing 20 potential numeric attributes like age and gender. Scientific Data (Spithourakis and Riedel, 2018) encompasses paragraphs extracted from Cornell’s ARXIV repository, featuring over half a million converted papers across 37 scientific sub-fields. Additionally, Elazar et al. (Elazar et al., 2019) proposed the Distributions over Quantities (DoQ) dataset, comprising empirical counts of scalar attribute values linked to more than 350K nouns, adjectives, and verbs across 10 different attributes.
@@
Furthermore, certain studies consider the relational knowledge between various numbers. For instance, the VERBPHYSICS (Forbes and Choi, 2017) dataset aggregates crowdsourced information about actions and objects, encompassing the relative knowledge of grounded object pairs and implications of actions associated with those objects. More intricately, the DROP (Dua et al., 2019) dataset necessitates Discrete Reasoning Over the content of Paragraphs. In this benchmark of 55,000 adversarially-created questions, a system is tasked with resolving references within a question, potentially spanning multiple input positions, and performing discrete operations like addition, counting, or sorting.
5.1.2. Math Word Problems
@@
A large number of datasets are proposed for MWPs. Hosseini et al. (Hosseini et al., 2014) curated the AddSub dataset, primarily focusing on addition and subtraction problems (Figure 9). This dataset identifies relevant variables and their respective values to translate sentences into problem statements, represented as equations. Another dataset, SingleOp (Roy et al., 2015), encompasses a wider range of mathematical operations including multiplication and division. Its purpose is to facilitate reasoning about quantities articulated in natural language. DRAW (Upadhyay and Chang, 2015) comprises 1000 algebra word problems semi-automatically annotated for evaluating automatic solvers. It features gold coefficient alignments crucial for uniquely identifying equation system derivations. The Alg514 (Kushman et al., 2014) dataset comprises 514 linear algebra problems structured around 28 equation templates. Meanwhile, MultiArith (Roy and Roth, 2015) is designed to handle arithmetic problems involving multiple steps and operations, without predefined templates. MATHPILE (Wang et al., 2023c) is a diverse and high-quality math-centric corpus comprising about 9.5 billion tokens. OpenWebMath (Paster et al., 2023) is an open dataset inspired by these works containing 14.7B tokens of mathematical webpages from Common Crawl.
@@
Several studies have introduced methods focused on predicting equations through semantic parsing. The SingleEq (Koncel-Kedziorski et al., 2015) dataset revolves around grade-school algebra word problems. It comprises samples that can be correlated to a single equation involving multiple mathematical operations on one variable, structured as a parsing tree (Figure 10). In contrast, the Dolphin1878 (Shi et al., 2015) dataset contains 1,878 word problems in mathematics, wherein each sample might encompass multiple equations and variables depicted through multiple trees. AllArith (Roy and Roth, 2017) offers a concise representation of dependencies among number units mentioned in a given problem, referred to as Unit Dependency Graphs (UDGs). To augment dataset size, Math23k (Wang et al., 2017) gathers 23,161 problems labeled with structured equations and corresponding answers. GSM8K (Cobbe et al., 2021) comprises 8.5K high-quality grade school math problems crafted by human problem writers. These problems typically involve 2 to 8 steps for resolution, predominantly requiring a sequence of elementary calculations using basic arithmetic operations to arrive at the final solution.
Figure 9. An example of AddSub, taken from (Hosseini et al., 2014).
Figure 10. An example of SingleEQ, taken from (Roy et al., 2015).
@@
HMWP (Qin et al., 2020) aims to enhance dataset diversity by extracting math word problems from a Chinese K12 problem bank. This effort sought to validate the universality of math word problem solvers and expand research in MWPs to better align with real-world scenarios. The dataset encompasses three types of MWPs: arithmetic word problems, equation set problems, and non-linear equation problems. In total, it comprises 5,491 MWPs, including 2,955 single-variable linear MWPs, 1,636 two-variable linear MWPs, and 900 single-variable nonlinear MWPs. Additionally, MathQA (Amini et al., 2019) presents a diverse collection of 37,000 English multiple-choice math word problems spanning various mathematical domains (Figure 11).
Figure 11. An example of MathQA, taken from (Amini et al., 2019).
@@
Existing research endeavors to enhance the comprehensibility and precision of intricate reasoning by providing fine-grained annotations of reasoning processes (Upadhyay and Chang, 2017; Ling et al., 2017; Chen et al., 2021; Lightman et al., 2023; Zhao et al., 2022). DRAW-1K (Upadhyay and Chang, 2017) introduces a novel dataset featuring 1,000 general algebra word problems, each annotated with derivations. AQUA (Ling et al., 2017) structures each question into four components: problem description, multiple-choice answer options, rationale, and correct option label (Figure 12). FinQA (Chen et al., 2021) presents an expert-annotated dataset comprising 8,281 financial QA pairs, meticulously annotated with numerical reasoning processes to ensure comprehensive explication. PRM800K (Lightman et al., 2023) assigns labels (positive, negative, or neutral) to each step in solving MATH problems sampled by a large-scale generator. The training set encompasses 800K step-level labels across 75K solutions to 12K problems.
Figure 12. An example of AQUA, taken from (Ling et al., 2017).
@@
Additionally, Patel et al. (Patel et al., 2021) introduced the SVAMP challenge set to establish a more robust evaluation framework for methods designed to tackle elementary-level math word problems. Alghamdi et al. (Alghamdi et al., 2022) contributed the inaugural large-scale dataset, ArMATH, for Arabic MWPs. This dataset comprises 6,000 samples of primary-school math problems written in Modern Standard Arabic. Kalyan et al. (Kalyan et al., 2021) proposed two datasets, REALFP and SYNTHFP, for a novel task called Fermi Problems (FPs). FPs entail questions whose answers can only be approximately estimated due to the impracticality or impossibility of precise computation. This task necessitates the amalgamation of various reasoning abilities, including suitable abstractions, commonsense knowledge, and the creative synthesis of problem-solving strategies.
@@
Recently, mathematical reasoning has incorporated multiple input modalities such as textual and tabular data (Lu et al., 2023b; Zhu et al., 2021; Zhao et al., 2022). For instance, TabMWP (Lu et al., 2023b) presents a new dataset comprising 38,431 open-domain grade-level problems that necessitate mathematical reasoning across both textual and tabular data. Zhu et al. (Zhu et al., 2021) introduced a large-scale QA dataset, TAT-QA, encompassing both tabular and textual data. This dataset often requires numerical reasoning—such as addition, subtraction, multiplication, division, counting, comparison/sorting, and their combinations—based on real financial reports to infer answers. Additionally, MultiHiertt (Zhao et al., 2022) is a proposed dataset consisting of QA pairs derived from Multi Hierarchical Tabular and Textual data (Figure 13). Each document in this dataset contains multiple tables and longer unstructured texts, accompanied by fine-grained annotations of reasoning processes and supporting facts to elucidate complex numerical reasoning.
Figure 13. An example of MultiHiertt, taken from (Zhao et al., 2022).
5.1.3. Theorem Proving